# CP violation in two-meson Tau decays induced by heavy new physics

### v1

@@ 4 Bounds on the New Physics coefficients
The Lagrangian eq. (2) is obtained as the relevant part of the low-energy limit of the SMEFT Lagrangian. This connection is useful to constrain the imaginary part of the scalar and tensor Wilson coefficients, as we explain in the following (see also refs. [39, 44]).
@@
$\mathcal{L}_{SMEFT}\supset[C^{(3)}_{\ell equ}]_{klmn}[(\bar{\nu}_{Lk}\sigma_{\mu\nu}e_{Rl})(\bar{d}_{Lm}\sigma^{\mu\nu}u_{Rn})-V_{am}(\bar{e}_{Lk}\sigma_{\mu\nu}e_{Rl})(\bar{u}_{La}\sigma^{\mu\nu}u_{Rn})]+\mathrm{h.c.}$ (24)
@@
The operators in the second term of eq. (24) contribute to the neutron electric dipole moment (through the $u$-quark contribution to it) and to $D^{0}-\bar{D}^{0}$ mixing, which constrain stringently $\Im m[\epsilon_{T}^{D}]$. We will consider, in our numerical analysis, two scenarios, as in ref. [44]:
  • If the neutron EDM  [^11] We recall that [44] $d_{n}=g_{T}^{u}(\mu)d_{u}(\mu)$, with $g_{T}^{u}(2\mathrm{\,GeV})=-0.204(14)$ [97, 98], $|d_{n}|<1.8\times 10^{-26}e$ cm at $90\%$ confidence level, C. L. [99]. is mainly contributed by a single $\hat{\epsilon}_{T}^{D}$, then
@@
    which yields $|\Im m[\hat{\epsilon}_{T}^{s}]|\lesssim 4\times 10^{-6}$ and $|\Im m[\hat{\epsilon}_{T}^{d}]|\lesssim 8\times 10^{-5}$ for $\Lambda\gtrsim 100$ GeV and $\mu=2$ GeV. The constraint on $D^{0}-\bar{D}^{0}$ mixing will then come from $\phi=\Im m\Big{[}V_{uD}V_{cD}\epsilon_{T(S)}^{D}\Big{]}$.
@@
We will consider $\phi=\pm\frac{\pi}{4}$, as in refs. [39, 44].
@@
Although the scalar operators do not enter directly the neutron EDM, they are still restricted indirectly because the bound on the tensor coefficients impacts the scalar ones through renormalization group equation evolution (details are given in ref. [44]). If we keep for simplicity only $C^{(1)}_{\ell equ}$ [44], $|\Im m[\hat{\epsilon}_{T}^{s}]|\lesssim 4\times 10^{-6}\Rightarrow|\Im m[\hat{\epsilon}_{S}^{s}]|\lesssim 2\times 10^{-3}$ and $|\Im m[\hat{\epsilon}_{T}^{d}]|\lesssim 8\times 10^{-5}\Rightarrow|\Im m[\hat{\epsilon}_{S}^{d}]|\lesssim 4\times 10^{-2}$. In the scalar case the strongest constraints come from $D^{0}-\bar{D}^{0}$ mixing, implying [44] $\Im m\left[\hat{\epsilon}^{D}_{S}\right]\in[-3.1,1.6]\times 10^{-4}$ at $95\%$ C.L., that is the restriction that dominates.
The real parts of the Wilson coefficients are most effectively constrained by analyzing CP-conserving inclusive and exclusive semi-leptonic tau decays [44, 50, 39, 76, 40, 85, 72, 100, 101, 102, 103]. We will take the results from ref. [103], which include improved radiative corrections (see also refs. [104, 105]), which are
$\Re e[\epsilon_{S}^{d}]=\left(1.0^{+0.6}_{-3.4}\right)\times 10^{-1}\,,\quad\Re e[\epsilon_{T}^{d}]=\left(0.4^{+4.3}_{-4.6}\right)\times 10^{-2}\,,$ (31)
with a correlation coefficient of $0.452$, and
@@ 5 Phenomenological analysis
Here we will present the numerical analysis of the most interesting observables sensitive to CP violation recalled in section 2, using the form factors introduced in section 3, and respecting the bounds on the Wilson coefficients of section 4. We will not discuss the $K_{S}\pi$ case, which has been studied extensively in e.g. refs. [44, 39, 40]  [^12] We have nevertheless reproduced fig. 3 in ref. [44], as this observable was not considered in ref. [40]. We thank Xin-Qian Li for useful feedback concerning this verification.. We will present the rate CP and the forward-backward asymmetries for the $\pi^{\pm}\pi^{0}$, $K^{\pm}K_{S}$ and $K^{\pm}\pi^{0}$ modes, in turn. Our results for $A_{CP}^{rate}$ will be quoted for the maximum possible signal, corresponding to $\Im m[\hat{\epsilon}_{T}^{d}]=\pm 8\times 10^{-5}$ and $\Im m[\hat{\epsilon}_{T}^{s}]=\pm 4\times 10^{-6}$ at $90\%$ C.L. In the case of $A_{FB}$ we will determine the optimal values of the $\hat{\epsilon}_{S,T}^{D}$ demanding that they maximize the following figure of merit  [^13] According to the preceding discussion we would naively expect that, in addition to the limits on $\Im m[\hat{\epsilon}_{S}^{D}]$, given at 90%C.L. by the interval $[-2.7,1.2]\times 10^{-4}$ –which determine $A_{CP}^{rate}$– the maximal CP violating signals would be found with the following Wilson coefficients (obtained from the extremes of the $90\%$C.L. intervals) $\Re e[\hat{\epsilon}_{S}^{d}]=-4.6\times 10^{-1}$, $\Re e[\hat{\epsilon}_{S}^{s}]=2.3\times 10^{-2}$ , for the scalar non-standard interactions, and $\Re e[\hat{\epsilon}_{T}^{d}]=-7.2\times 10^{-2}$ and $\Re e[\hat{\epsilon}_{T}^{s}]=1.8\times 10^{-2}$, for the tensor interactions.
$F(\hat{\epsilon}_{D}^{D},\hat{\epsilon}_{T}^{D})=\int_{(m_{1}+m_{2})^{2}}^{M_{\tau}^{2}}\Big{|}A_{FB}(s;\hat{\epsilon}_{S}^{D},\hat{\epsilon}_{T}^{D})\Big{|}\quad ds\,.$ (33)
@@ 5.1 CP violating asymmetries in the $\pi^{\pm}\pi^{0}$ channel
In this case the SM contribution is negligible, so using eq. (11) for the NP part, we find
@@
The maximal $A_{FB}|_{\pi\pi}$ is shown by a blue solid line in figure 5, which was obtained with the values $\Re e[\epsilon_{S}^{d}]=-0.46$ , $\Im m[\epsilon_{S}^{d}]=-2.7\times 10^{-4}$, $\Re e[\epsilon_{T}^{d}]=-7.2\times 10^{-2}$ and $\Im m[\epsilon_{T}^{d}]=8\times 10^{-5}$. For comparison, the SM contribution (very close to zero, except at threshold) is shown by the black dashed line. Since $|A_{FB}(s)|\geq 0.2$ for most of the spectrum, it would be feasible to measure such asymmetries at Belle-II and future experiments.
Figure 5: Maximal $A^{\tau\rightarrow\pi\pi^{0}\nu_{\tau}}_{FB}(s)$ (blue solid line), corresponding to the Wilson coefficients values $\Re e[\epsilon_{S}^{d}]=-0.46$ , $\Im m[\epsilon_{S}^{d}]=-2.7\times 10^{-4}$, $\Re e[\epsilon_{T}^{d}]=-7.2\times 10^{-2}$ and $\Im m[\epsilon_{T}^{d}]=8\times 10^{-5}$, compared to the SM case (black dashed line).
In the SM, the FoM $F(\hat{\epsilon}_{S}=0,\hat{\epsilon}_{T}=0)$ reads $7\times 10^{-3}$. Allowing the Wilson coefficients to vary within their allowed intervals, we find that its maximum value is $0.64$, which is obtained for $\Re e[\epsilon_{S}^{d}]=-0.46$, $\Im m[\epsilon_{S}^{d}]=1.7\times 10^{-4}$, $\Re e[\epsilon_{T}^{d}]=-7.1\times 10^{-2}$ and $\Im m[\epsilon_{T}^{d}]=-8\times 10^{-5}$. For this maximal NP allowed, measurement shall be possible at present experiments. Fig. 6 shows a three-dimensional density plot of the FoM $F$ for fixed $\Im m[\hat{\epsilon}_{T}^{d}]$, varying its real part and the complex number $\hat{\epsilon}_{S}^{d}$.
@@ 5.2 CP violating asymmetries in the $K^{\pm}K_{S}$ channel
As we will see, this is the most interesting channel. Using eq. (12) we find
$A_{CP}^{rate}|_{KK}=3.8\times 10^{-3},$ (35)
@@
that is dominated by the SM contribution (for which we used the result adapted to the BaBar setting, with a SM CP rate asymmetry of $3.6\times 10^{-3}$) and
$A_{CP}^{rate}|_{KK,NP}\leq 2.3\times 10^{-4},$ (36)
@@
obtained with the analog for this channel of eq. (11). Noticeably, a $5\%$ precision on the measurement of $A_{CP}^{rate}$ would already be sensitive to the maximum allowed NP contribution in this case. This contrasts sharply with the $K_{S}\pi^{\pm}$ channels, where the NP contributions is orders of magnitude smaller than the SM one, complicating the explanation of the BaBar anomaly. There are two reasons why this NP contribution can be so much larger than for the $K_{S}\pi^{\pm}$ modes. On the one hand, the bound on the Wilson coefficient is a factor twenty larger here ($|\Im m[\hat{\epsilon}_{T}^{d}]|\leq 8\times 10^{-5}$ vs. $|\Im m[\hat{\epsilon}_{T}^{s}]|\leq 4\times 10^{-6}$ [^14] We recall that $A_{CP}^{rate}$ is linear on $\Im m[\hat{\epsilon}_{T}^{D}]$, eq.(11).. On the other, the $K_{S}K^{\pm}$ channel lies in the inelastic region, while the main contribution to the decay width of the $K_{S}\pi^{\pm}$ belongs to the elastic regime, where $A_{CP}^{rate}$ vanishes identically. Due to these facts, we get such a large NP effect, that can be up to $\mathcal{O}(5\%)$ on $A_{CP}^{rate}|_{KK}$. We therefore encourage the experimental collaborations to measure this observable trying to reach this level of precision. On the other hand, a measurement with $10\%$ accuracy of $A_{CP}^{rate}$ in the $K_{S}K$ channels should coincide with the result for $K_{S}\pi$, so it will be a purely experimental test of the BaBar anomaly that we prompt BaBar and Belle(-II) to perform.
@@
The maximal $A_{FB}|_{KK}$ is shown by a blue solid line in figure 7, which is obtained with the same set of Wilson coefficients as in the previous subsection, $\Re e[\epsilon_{S}^{d}]=-0.46$ , $\Im m[\epsilon_{S}^{d}]=-2.7\times 10^{-4}$, $\Re e[\epsilon_{T}^{d}]=-7.2\times 10^{-2}$ and $\Im m[\epsilon_{T}^{d}]=8\times 10^{-5}$. The (unmeasurably small) SM contribution is represented by a black dashed line. The obtained asymmetries, of the order of $-0.5$ in a big portion of the spectrum, should be observable at current and future facilities.
@@
Figure 7: Maximal $A^{\tau\rightarrow K_{S}K\nu_{\tau}}_{FB}(s)$ (blue solid line), corresponding to the Wilson coefficients values $\Re e[\epsilon_{S}^{d}]=-0.46$ , $\Im m[\epsilon_{S}^{d}]=-2.7\times 10^{-4}$, $\Re e[\epsilon_{T}^{d}]=-7.2\times 10^{-2}$ and $\Im m[\epsilon_{T}^{d}]=8\times 10^{-5}$, compared to the SM case (black dashed line).
@@
In the SM, the FoM $F(\hat{\epsilon}_{S}=0,\hat{\epsilon}_{T}=0)$ reads $1.8\times 10^{-2}$. Varying the Wilson coefficients within their allowed ranges, we find that its maximum value is $0.82$, which is obtained for $\Re e[\epsilon_{S}^{d}]=-0.46$, $\Im m[\epsilon_{S}^{d}]=-2.7\times 10^{-4}$, $\Re e[\epsilon_{T}^{d}]=-7.1\times 10^{-2}$ and $\Im m[\epsilon_{T}^{d}]=8\times 10^{-5}$. For this maximal NP allowed, measurement shall be possible at present experiments. Fig. 8 shows a three-dimensional density plot of the FoM $F$ for fixed $\Im m[\hat{\epsilon}_{T}^{d}]$, varying its real part and the complex number $\hat{\epsilon}_{S}^{d}$.

### v2

@@ 4 Bounds on the New Physics coefficients
The Lagrangian eq. (2) is obtained as the relevant part of the low-energy limit of the SMEFT Lagrangian, integrating out the heavy particles, that are non-dynamical at the considered scales. This connection is useful to constrain the imaginary part of the scalar and tensor Wilson coefficients, as we explain in the following (see also refs. Cirigliano:2017tqn ; Chen:2021udz ).
@@
$\mathcal{L}_{SMEFT}$ $\supset[C^{(3)}_{\ell equ}]_{klmn}[(\bar{\nu}_{Lk}\sigma_{\mu\nu}e_{Rl})(\bar{d}_{Lm}\sigma^{\mu\nu}u_{Rn})-V_{am}(\bar{e}_{Lk}\sigma_{\mu\nu}e_{Rl})(\bar{u}_{La}\sigma^{\mu\nu}u_{Rn})]$ (24)
$\quad+\mathrm{h.c.}$
@@
The operators in the second term of eq. (24) contribute to the neutron electric dipole moment (through the $u$-quark contribution to it) and to $D^{0}-\bar{D}^{0}$ mixing, which constrain stringently $\Im m[\epsilon_{T}^{D}]$. We will consider, in our numerical analysis, two scenarios, as in ref. Chen:2021udz :
  • If the neutron EDM  [^11] We recall that Chen:2021udz $d_{n}=g_{T}^{u}(\mu)d_{u}(\mu)$, with $g_{T}^{u}(2\mathrm{\,GeV})=-0.204(14)$ Bhattacharya:2015esa ; Gupta:2018lvp , $|d_{n}|<1.8\times 10^{-26}e$ cm at $90\%$ confidence level, C. L. Abel:2020pzs . is mainly contributed by a single $\hat{\epsilon}_{T}^{D}$, then
@@
    which yields $|\Im m[\hat{\epsilon}_{T}^{s}]|\lesssim 4\times 10^{-6}$ and $|\Im m[\hat{\epsilon}_{T}^{d}]|\lesssim 8\times 10^{-5}$ for $\Lambda\gtrsim 100$ GeV and $\mu=2$ GeV  [^12] The relative strength of these bounds is explained by the small $|V_{us}/V_{ud}|^{2}$ ratio.. The constraint on $D^{0}-\bar{D}^{0}$ mixing will then come from $\phi=\Im m\Big{[}V_{uD}V_{cD}\epsilon_{T(S)}^{D}\Big{]}$.
@@
We will consider $\phi=\pm\frac{\pi}{4}$, as in refs. Cirigliano:2017tqn ; Chen:2021udz .
@@
Although the scalar operators do not enter directly the neutron EDM, they are still restricted indirectly because the bound on the tensor coefficients impacts the scalar ones through renormalization group equation evolution (details are given in ref. Chen:2021udz ). If we keep for simplicity only $C^{(1)}_{\ell equ}$ Chen:2021udz , $|\Im m[\hat{\epsilon}_{T}^{s}]|\lesssim 4\times 10^{-6}\Rightarrow|\Im m[\hat{\epsilon}_{S}^{s}]|\lesssim 2\times 10^{-3}$ and $|\Im m[\hat{\epsilon}_{T}^{d}]|\lesssim 8\times 10^{-5}\Rightarrow|\Im m[\hat{\epsilon}_{S}^{d}]|\lesssim 4\times 10^{-2}$. In the scalar case the strongest constraints come from $D^{0}-\bar{D}^{0}$ mixing, implying Chen:2021udz $\Im m\left[\hat{\epsilon}^{D}_{S}\right]\in[-3.1,1.6]\times 10^{-4}$ at $95\%$ C.L., that is the restriction that dominates.
The real parts of the Wilson coefficients $\epsilon_{T(S)}^{s}$ are most effectively constrained by analyzing CP-conserving inclusive and exclusive semi-leptonic tau decays Chen:2021udz ; Garces:2017jpz ; Cirigliano:2017tqn ; Miranda:2018cpf ; Rendon:2019awg ; Gonzalez-Solis:2019lze ; Gonzalez-Solis:2020jlh ; Cirigliano:2018dyk ; Cirigliano:2021yto ; Arteaga:2022xxy ; Escribano:2023seb . For the strangeness-changing sector, we will take the results from ref. Escribano:2023seb that include improved radiative corrections (see also refs. Arroyo-Urena:2021nil ; Arroyo-Urena:2021dfe ), which are
@@
The real parts of the Wilson coefficients $\epsilon_{T(S)}^{d}$, entering strangeness-conserving processes, were better constrained in Cirigliano:2018dyk
$\Re e[\epsilon_{S}^{d}]=\left(-0.6\pm 1.5\right)\times 10^{-2}\,,\quad\Re e[\epsilon_{T}^{d}]=\left(-0.04\pm 0.46\right)\times 10^{-2}\,,$ (32)
with a null correlation coefficient.
@@ 5 Phenomenological analysis
Here we will present the numerical analysis of the most interesting observables sensitive to CP violation recalled in section 2, using the form factors introduced in section 3, and respecting the bounds on the Wilson coefficients of section 4. We will not discuss the $K_{S}\pi$ case, which has been studied extensively in e.g. refs. Chen:2021udz ; Cirigliano:2017tqn ; Rendon:2019awg  [^13] We have nevertheless reproduced figure 3 in ref. Chen:2021udz , as this observable was not considered in ref. Rendon:2019awg . We thank Xin-Qian Li for useful feedback concerning this verification. We will present the rate CP and the forward-backward asymmetries for the $\pi^{\pm}\pi^{0}$, $K^{\pm}K_{S}$ and $K^{\pm}\pi^{0}$ modes, in turn. Our results for $A_{CP}^{rate}$ will be quoted for the maximum possible signal, corresponding to $\Im m[\hat{\epsilon}_{T}^{d}]=\pm 8\times 10^{-5}$ and $\Im m[\hat{\epsilon}_{T}^{s}]=\pm 4\times 10^{-6}$ at $90\%$ C.L. In the case of $A_{FB}$ we will determine the optimal values of the $\hat{\epsilon}_{S,T}^{D}$ demanding that they maximize the following figure of merit  [^14] According to the preceding discussion we would naively expect that, in addition to the limits on $\Im m[\hat{\epsilon}_{S}^{D}]$, given at 90%C.L. by the interval $[-2.7,1.2]\times 10^{-4}$ –which determine $A_{CP}^{rate}$– the maximal CP violating signals would be found with the following Wilson coefficients (obtained from the extremes of the $90\%$C.L. intervals) $\Re e[\hat{\epsilon}_{S}^{d}]=-3.1\times 10^{-2}$, $\Re e[\hat{\epsilon}_{S}^{s}]=2.3\times 10^{-2}$ , for the scalar non-standard interactions, and $\Re e[\hat{\epsilon}_{T}^{d}]=-7.9\times 10^{-3}$ and $\Re e[\hat{\epsilon}_{T}^{s}]=1.8\times 10^{-2}$, for the tensor interactions.
$F(\hat{\epsilon}_{S}^{D},\hat{\epsilon}_{T}^{D})=\int_{(m_{1}+m_{2})^{2}}^{M_{\tau}^{2}}\Big{|}A_{FB}(s;\hat{\epsilon}_{S}^{D},\hat{\epsilon}_{T}^{D})\Big{|}\quad ds\,.$ (33)
@@ 5.1 CP violating asymmetries in the $\pi^{\pm}\pi^{0}$ channel
In this case the SM contribution is negligible, so using eq. (11) for the NP part, we find that
@@
The maximal $A_{FB}|_{\pi\pi}$ is shown by a blue solid line in figure 5, which was obtained with the values $\Re e[\epsilon_{S}^{d}]=-3.1\times 10^{-2}$ , $\Im m[\epsilon_{S}^{d}]=-2.7\times 10^{-4}$, $\Re e[\epsilon_{T}^{d}]=-7.9\times 10^{-3}$ and $\Im m[\epsilon_{T}^{d}]=8\times 10^{-5}$. For comparison, the SM contribution (very close to zero, except at threshold) is shown by the black dashed line. Since $|A_{FB}(s)|\leq 0.1$ for most of the spectrum, it will be challenging to measure such asymmetries at Belle-II and future experiments.
Figure 5: Maximal $A^{\tau\rightarrow\pi\pi^{0}\nu_{\tau}}_{FB}(s)$ (blue solid line), corresponding to the Wilson coefficients values $\Re e[\epsilon_{S}^{d}]=-3.1\times 10^{-2}$ , $\Im m[\epsilon_{S}^{d}]=-2.7\times 10^{-4}$, $\Re e[\epsilon_{T}^{d}]=-7.9\times 10^{-3}$ and $\Im m[\epsilon_{T}^{d}]=8\times 10^{-5}$, compared to the SM case (black dashed line).
In the SM, the FoM $F(\hat{\epsilon}_{S}=0,\hat{\epsilon}_{T}=0)$ reads $1.2\times 10^{-2}$. Allowing the Wilson coefficients to vary within their allowed intervals, we find that its maximum value is $4.1\times 10^{-2}$ , which is obtained for $\Re e[\epsilon_{S}^{d}]=-3.1\times 10^{-2}$, $\Im m[\epsilon_{S}^{d}]=1.7\times 10^{-4}$, $\Re e[\epsilon_{T}^{d}]=-7.9\times 10^{-3}$ and $\Im m[\epsilon_{T}^{d}]=-8\times 10^{-5}$. Even for this maximal NP allowed, measurement will be challenging. Figure 6 shows a three-dimensional density plot of the FoM $F$ for fixed $\Im m[\hat{\epsilon}_{T}^{d}]$, varying its real part and the complex number $\hat{\epsilon}_{S}^{d}$.
@@ 5.2 CP violating asymmetries in the $K^{\pm}K_{S}$ channel
As we will see, this is the most interesting channel. Using eq. (12) and the negative bound for $\Im m[\epsilon_{T}^{d}]$, allows us to get the maximum absolute value for $A_{CP}^{rate}|_{KK}$, in order to obtain  [^15] We note that, for the $K_{S}K^{\pm}$ channel, the expected standard model contribution $A_{CP}^{rate}|_{KK,SM}$ is opposite in sign to that of the channel $K_{S}\pi^{\pm}$.
$A_{CP}^{rate}|_{KK}=-3.83\times 10^{-3},$ (35)
@@
that is dominated by the SM contribution (for which we used the result adapted to the BaBar setting, with a SM CP rate asymmetry of $-3.6\times 10^{-3}$) and
$|A_{CP}^{rate}|_{KK,NP}|\leq 2.3\times 10^{-4},$ (36)
found with the analog for this channel of eq. (11). Actually, by considering the interval in (36), we can set lower and upper bounds for $A_{CP}^{rate}|_{KK}$, i.e.
$-3.83\times 10^{-3}\leq A_{CP}^{rate}|_{KK}\leq-3.37\times 10^{-3}.$ (37)
@@
Noticeably, a $5\%$ precision on the measurement of $A_{CP}^{rate}$ would already be sensitive to the maximum allowed NP contribution in this case. This contrasts sharply with the $K_{S}\pi^{\pm}$ channels, where the NP contributions is orders of magnitude smaller than the SM one, complicating the explanation of the BaBar anomaly. There are two reasons why this NP contribution can be so much larger than for the $K_{S}\pi^{\pm}$ modes. On the one hand, the bound on the Wilson coefficient is a factor twenty larger here ($|\Im m[\hat{\epsilon}_{T}^{d}]|\leq 8\times 10^{-5}$ vs. $|\Im m[\hat{\epsilon}_{T}^{s}]|\leq 4\times 10^{-6}$).  [^16] We recall that $A_{CP}^{rate}$ is linear on $\Im m[\hat{\epsilon}_{T}^{D}]$, eq. (11). On the other, the $K_{S}K^{\pm}$ channel lies in the inelastic region, while the main contribution to the decay width of the $K_{S}\pi^{\pm}$ belongs to the elastic regime, where $A_{CP}^{rate}$ vanishes identically. Due to these facts, we get such a large NP effect, that can be up to $\mathcal{O}(5\%)$ on $A_{CP}^{rate}|_{KK}$. We therefore encourage the experimental collaborations to measure this observable trying to reach this level of precision. On the other hand, a measurement with $10\%$ accuracy of $A_{CP}^{rate}$ in the $K_{S}K$ channels should coincide (up to sign) with the result for $K_{S}\pi$, so it will be a purely experimental test of the BaBar anomaly that we prompt BaBar and Belle(-II) to perform. As a reference, in the BaBar measurement BaBar:2018qry , the uncertainty of the best measured points is $\sim 3\%$, so these numbers are not unrealistic targets.
@@
The maximal $A_{FB}|_{KK}$ is shown by a blue solid line in figure 7, which is obtained with the set of Wilson coefficients, $\Re e[\epsilon_{S}^{d}]=-3.1\times 10^{-2}$ , $\Im m[\epsilon_{S}^{d}]=-2.7\times 10^{-4}$, $\Re e[\epsilon_{T}^{d}]=-7.1\times 10^{-3}$ and $\Im m[\epsilon_{T}^{d}]=8\times 10^{-5}$. The (unmeasurably small) SM contribution is represented by a black dashed line. The obtained asymmetries, of the order of $-0.1$ in a big portion of the spectrum, should be observable at current and future facilities.
@@
Figure 7: Maximal $A^{\tau\rightarrow K_{S}K\nu_{\tau}}_{FB}(s)$ (blue solid line), corresponding to the Wilson coefficients values $\Re e[\epsilon_{S}^{d}]=-3.1\times 10^{-2}$ , $\Im m[\epsilon_{S}^{d}]=-2.7\times 10^{-4}$, $\Re e[\epsilon_{T}^{d}]=7.1\times 10^{-3}$ and $\Im m[\epsilon_{T}^{d}]=8\times 10^{-5}$, compared to the SM case (black dashed line).
@@
In the SM, the FoM $F(\hat{\epsilon}_{S}=0,\hat{\epsilon}_{T}=0)$ reads $1.1\times 10^{-2}$. Varying the Wilson coefficients within their allowed ranges, we find that its maximum value is $0.14$, which is obtained for $\Re e[\epsilon_{S}^{d}]=-3.1\times 10^{-2}$, $\Im m[\epsilon_{S}^{d}]=-2.7\times 10^{-4}$, $\Re e[\epsilon_{T}^{d}]=7.1\times 10^{-3}$ and $\Im m[\epsilon_{T}^{d}]=8\times 10^{-5}$. For this maximal NP allowed, measurement shall be possible at present experiments. Figure 8 shows a three-dimensional density plot of the FoM $F$ for fixed $\Im m[\hat{\epsilon}_{T}^{d}]$, varying its real part and the complex number $\hat{\epsilon}_{S}^{d}$.